# LAVA: Explainability for Unsupervised Latent Embeddings

### v1

@@ Abstract
Unsupervised black-box models can be drivers of scientific discovery, but remain difficult to interpret. Crucially, discovery hinges on understanding the model output, which is often a multi-dimensional latent embedding rather than a well-defined target. While explainability for supervised learning usually seeks to uncover how input features are used to predict a target, its unsupervised counterpart should relate input features to the structure of the learned latent space. Adaptations of supervised model explainability for unsupervised learning provide either single-sample or dataset-wide summary explanations. However, without automated strategies of relating similar samples to one another guided by their latent proximity, explanations remain either too fine-grained or too reductive to be meaningful. This is especially relevant for manifold learning methods that produce no mapping function, leaving us only with the relative spatial organization of their embeddings. We introduce Locality-Aware Variable Associations (LAVA), a post-hoc model-agnostic method designed to explain local embedding organization through its relationship with the input features. To achieve this, LAVA represents the latent space as a series of localities (neighborhoods) described in terms of correlations between the original features, and then reveals reoccurring patterns of correlations across the entire latent space. Based on UMAP embeddings of MNIST and a single-cell kidney dataset, we show that LAVA captures relevant feature associations, with visually and biologically relevant local patterns shared among seemingly distant regions of the latent spaces.
@@ 1 Introduction
A pragmatic explanation can simply be defined as an answer to the question “why?” [1]. Traditional models, like Newton’s law of gravitation, constitute pragmatic explanations as they allow reasoning about observable phenomena through rigorous mathematical description. Similarly, stochastic phenomena that lend themselves well to statistical modeling, including linear regression, are explainable through probabilistic reasoning. For more complex problems such as facial recognition or protein structure prediction, expert-led modeling is superseded by machine learning (ML), where flexible model architectures and automated end-to-end optimization algorithms take center stage. ML has seen considerable progress and success in leveraging high-dimensional data for scientific discovery across fields like chemistry, mathematics, and biology [2, 3, 4, 5]. However, even when ML models make reliable predictions or reflect observed phenomena, reasoning about their decisions is notoriously difficult as many ML models are “black boxes”: even with full decomposibility, the complexity of the data and operations required to produce an output often exceeds what is practically comprehensible [6, 7, 8]. Limited insight into model behavior is a major barrier to wider adoption of ML for decision making, especially in critical areas such as healthcare and aviation [9, 10, 5, 11]. To improve the understanding of ML, the community uses either intrinsically interpretable but restrictive models, or post-hoc explainability methods that enable interpretation of black-box models [9, 12, 13, 10, 14].
@@
The increasing popularity and complexity of black-box ML poses challenges to existing post-hoc explainability methods, exposing their shortcomings and creating a need for new solutions. At its core, post-hoc explainability aims to enable reasoning about the output of the model in terms of its treatment of the original input data. This information enables cross-referencing of model behavior with domain knowledge for further verification and assessment of model reliability, as well as mitigation of unwanted biases. Existing explainability methods cover a wide range of approaches, including feature or sample importance, anchors, counterfactual explanations, among others [14]. However, most have been developed to explain supervised learning models, whose output tends to be a well-defined target label or quantity. This ostensibly ignores unsupervised black-box learning strategies, which are designed to reveal latent structure with limited prior knowledge, and thus also contribute meaningfully to scientific discovery. Indeed, latent embeddings based on autoencoders, transformers, and diffusion models have shown effectiveness in various tasks [15, 16, 17, 18], and serve as the basis of foundation models [19]. Similarly, manifold learning is prominently used in bioinformatics for data exploration, trajectory inference, cell type annotation, spatial domain identification, and multi-omics integration [18, 20, 21, 22, 23].
@@
Crucially, most unsupervised black-box models output multi-dimensional latent embeddings. Relating such dimensions to the input data is less straightforward, as the axes of the learned latent space do not correspond to known quantities and are therefore difficult to interpret [24]. Approaches for disentangled representations have shown mixed results, suggesting that latent dimensions cannot be explained independently from one another [24, 25]. As a result, it might be necessary for explainability of unsupervised black-box models to focus on relationships between input data and the relative spatial organization of the embeddings. Moreover, many manifold learning methods only produce representations of the samples in the learned latent space, without an underlying mapping function to embed new data [26]. This means that all information is contained exclusively within the embeddings, making the link between latent organization and original data crucial for reasoning about the underlying model.
@@
Existing literature on explainability of unsupervised black-box latent embeddings is limited. Most notably, [24] introduce a framework adapting feature and sample importance methodology developed for supervised models to the unsupervised setting. This contribution importantly enables explanations for unsupervised models using a wide range of explainability methods. However, available methods yield either individual sample or dataset-wide summary explanations, which ignore the inherent organization of the latent embeddings and are likely too fine-grained or too global to meaningfully understand the underlying model. This is especially relevant as many black-box models are non-linear and seek to capture local structure in the original high-dimensional data. In fact, the popular t-SNE and UMAP manifold learning algorithms  [27, 28, 29] explicitly aim to preserve local sample neighborhoods of the input space in a lower-dimensional latent space. We therefore argue that effective explainability of latent embeddings requires some strategy to relate samples to one another based on their organization or proximity in the latent space. Manifold learning methods specifically have been characterized in terms of neighborhood retention [30], and further analyzed to recommend ways of setting hyperparameters [31, 18, 32, 33]. While these studies offer insight into these methods, they do not provide information on how the embeddings relate to the the original data. More manual approaches can relate spatial organization to original features [34], but rely on prior knowledge, metadata, or target features of interest, and are likely infeasible for higher-dimensional data.
In this work, we propose Locality-Aware Variable Associations (LAVA), a model-agnostic method to explain learned latent embeddings through relationships between the input features and the local latent spatial organization. To explain local structure, LAVA creates explanations based on localities: neighborhoods of sample embeddings across the latent space. Each latent locality is linked to the original data through a representation comprising all pairwise correlations between the original features, considering only the original samples belonging to that locality’s neighborhood. To reveal the consistency of local relationships between localities throughout the latent space, LAVA identifies modules of reoccurring subpatterns of feature correlations across localities. Experiments with LAVA reveal that neighborhoods of embeddings across a latent space can show distinct correlation patterns but also share subpatterns of correlation, and demonstrate that LAVA modules capture similar subpatterns between localities. We use LAVA to explain UMAP embeddings of both MNIST and single-cell kidney datasets, and showcase its ability to identify visually and biologically relevant modules of pixels and gene expression correlation, respectively.
@@ 2 Methodology
The LAVA method consists of three steps: (1) locality definition, (2) locality representation, and (3) module extraction (Figure 1). The first two steps create a reduced representation of the embeddings to explain, linking the latent and original spaces through local latent neighborhoods called localities, summarized by correlations of the original features. The third step uncovers structure preserved throughout the latent space, in the form of correlation subpatterns reoccurring across localities.
Figure 1: LAVA method overview. a, LAVA explains latent embeddings of a dataset (example UMAP embeddings of MNIST). b, Based on a neighborhood size hyperparameter $k$, LAVA represents the latent space as a series of localities: overlapping neighborhoods centered around latent points called probes. c, For each locality, LAVA calculates feature-to-feature correlations using the original samples of that locality (top right, locality correlations; bottom right, heatmap showing the sum of correlations per pixel, with added lines connecting correlated pixels). d, LAVA identifies modules yielding shared subpatterns of correlation across localities.
@@ 2.1 Locality Definition
To reflect the local structure of the input latent sample embeddings, LAVA uses a smaller number of localities centered around probe points (not samples) that are spread across the latent space (Figure 1b). The locality centered at probe $p$ is then the $n$-sized Euclidean neighborhood containing the $n$ neighbor samples nearest to $p$ in the latent space. LAVA defines the number of localities $\ell\leq E$ as $\ell=E\times o/n$, based on the number of input embeddings $E$ and two hyperparameters: the neighborhood size $n$, and a factor $o$ controlling the overlap between localities (Appendix A.1).
Locality placement. LAVA determines the placement of localities by optimizing locality in-degree centrality towards neighborhood in-degree centrality. In other words, by making the number of probe-based localities including a sample as a neighbor approximate, in relative terms, the number of $n$-neighborhoods including a sample as a neighbor out of all sample-based $n$-sized neighborhoods (Figures S5 and S11). Locality placement is optimized using weighted $k$-means clustering in combination with the DIRECT optimization algorithm [35, 36]: the goal is to find the locations of the $\ell$ probes as cluster centers in $k$-means, with $k=\ell$, around which the $\ell$ localities will be formed. Clustering is performed using the following sample weighting scheme:
$\centering w_{i}=\left(\frac{in\_neighborhood_{i}}{E}\right)^{\alpha}\times\left(\frac{1}{\bar{d}_{i}^{n}}\right)^{\beta},\@add@centering$ (1)
where $w_{i}$ is the weight of sample $i$, $in\_neighborhood_{i}$ is the in-degree centrality of sample $i$ using sample-based $n$-neighborhoods, $\bar{d}_{i}^{n}$ is the average distance from sample $i$ to its $n$ nearest neighbors, and $\alpha$ and $\beta$ are parameters optimized using the DIRECT algorithm (Appendix A.1). We define the cost function for DIRECT as the Manhattan distance between the neighborhood and locality in-degree centralities of each sample, as follows:
$\centering\mathcal{L}(\alpha,\beta)=\sum_{i=1}^{E}\left|\frac{in\_neighborhood_{i}}{E}-\frac{in\_locality_{i,\alpha,\beta}}{\ell}\right|,\@add@centering$ (2)
where $in\_locality_{i,\alpha,\beta}$ is the number of localities around the resulting $\ell$ cluster centers that include $i$ as a neighbor. Implementation details and pseudocode can be found in Appendix A.1.
@@ 2.2 Locality Representation
After defining localities, LAVA links the local structure of the latent space to the original data. This is achieved by representing each locality through pairwise associations between features, using only the samples within that locality in their original feature space. In this version of LAVA, feature association is defined as the absolute value of the Spearman correlation coefficient, capturing the strength of monotone association between features as values in $[0,1]$ (Figure 1c). Hence, the representation of each locality is a vector of $D\choose 2$ pairwise correlations for an original dataset with $D$ features. Unlike Pearson correlations, Spearman correlations are not scale dependent and thus more flexible in the type of association they capture, while remaining computationally inexpensive and interpretable downstream. Individual correlations are filtered (set to 0) if one of the original features has the same value in $75\%$ of the neighborhood (Appendix A.2).
@@ 2.3 Module Extraction
Finally, to reveal local relations preserved throughout the latent space, LAVA identifies modules of reoccurring feature association subpatterns across localities (Figure 2). Modules adhere to two properties: (1) modules capture reoccurring rather than average patterns across localities, since modules based on averages of localities would not be accurate representations of correlations observed at each locality individually; (2) modules can be co-present in the same locality, allowing decomposition of the locality pattern of correlation into subpatterns.
Extraction. Modules are identified based on the matrix of locality representations $\boldsymbol{C}$, with rows denoting the $\ell$ localities and columns representing the $D\choose 2$} pairwise correlations. To identify modules, we propose the Association Matrix Factorization (AMF) method, which factorizes a matrix like $\boldsymbol{C}$ into two: module matrix $\boldsymbol{M}$ and presence matrix $\boldsymbol{P}$. The module matrix $\boldsymbol{M}$ is of size $M\times{D\choose 2}$, where each row represents one of $M$ modules, columns denote correlations, and each row vector expresses a subpattern of correlations for the corresponding module. The presence matrix $\boldsymbol{P}$ is of size $\ell\times M$, with each entry quantifying the presence of one of the $M$ modules (or its subpattern) in one of the $\ell$ localities. Both presence and module matrix entries are restricted to real values in $[0,1]$, and they reconstruct the locality correlations using the maximum formulation described below.
Figure 2: Module-based factorization of localities. a, A locality $i$ is reconstructed by multiplying each module’s entries by the module presence for that locality, and then taking the maximum per entry across the $M$ extracted modules. b, Localities are represented as a matrix $\boldsymbol{C}$, where each row is a locality, each column a pair of original features, and each entry the pairwise feature correlation in that locality. Module extraction approximates $\boldsymbol{C}$ using a maximum per entry based on modules $\boldsymbol{M}$ and their presences $\boldsymbol{P}$: it calculates an outer product between each column of $\boldsymbol{P}$ (superscript) and each row of $\boldsymbol{M}$, yielding a tensor of $M$ matrix slices of the same shape as $\boldsymbol{C}$ (Equation 5). Each slice represents a module’s contribution to the correlations of a locality. The $M$ dimension is reduced by taking the per element maximum, yielding the reconstructed locality dataset.
2.3.1 AMF formulation
The formulation of AMF enables modules to overlap across localities. Crucially, as correlations or measures of associations more broadly cannot be considered additive, we avoid summing correlations and instead use a per element maximum function. This allows modules to contribute potentially overlapping subpatterns of correlations without direct interaction in the reconstruction. To that effect, we reconstruct each correlation $j$ of each locality $i$ by taking the maximum of the modules’ corresponding correlation values multiplied by the presences of those module at locality $i$:
@@ 2.3.1 AMF formulation
where $\odot$ is the Hadamard product and single letter subscripts indicate matrix rows. The reconstruction of the entire locality $i$ can then be written as a per element maximum across products of the locality’s presence values and the modules (Figure 2a):
$\centering\boldsymbol{\hat{C}}_{i}=\max\limits_{1\leq j\leq{D\choose 2}}\left(\boldsymbol{P}_{i1}\times\boldsymbol{M}_{1}^{T},\boldsymbol{P}_{i2}\times\boldsymbol{M}_{2}^{T},...,\boldsymbol{P}_{iM}\times\boldsymbol{M}_{M}^{T}\right),\@add@centering$ (4)
@@
$\centering\boldsymbol{P}^{m}\otimes\boldsymbol{M}_{m}^{T}=\begin{bmatrix}\boldsymbol{P}_{1m}\boldsymbol{M}_{m1}&\boldsymbol{P}_{1m}\boldsymbol{M}_{m2}&\ldots&\boldsymbol{P}_{1m}\boldsymbol{M}_{m{D\choose 2}}\\ \boldsymbol{P}_{2m}\boldsymbol{M}_{m1}&\boldsymbol{P}_{2m}\boldsymbol{M}_{m2}&\ldots&\boldsymbol{P}_{2m}\boldsymbol{M}_{m{D\choose 2}}\\ \vdots&\vdots&\ddots&\vdots\\ \boldsymbol{P}_{\ell m}\boldsymbol{M}_{m1}&\boldsymbol{P}_{\ell m}\boldsymbol{M}_{m2}&\ldots&\boldsymbol{P}_{\ell m}\boldsymbol{M}_{m{D\choose 2}}\end{bmatrix}$,\@add@centering (5)
where the superscript $m$ denotes the column of the matrix $P$ (the presence vector of some module $m$) and $\otimes$ is the outer product. Intuitively, the reconstruction of all localities is similar to matrix multiplication between the presence matrix and the module matrix, but instead of summing up entries across the module dimension, we take their maximum. In tensor terminology, outer products from Equation 5 for each of the modules produce $M$ slices of size $\ell\times{D\choose 2}$, over which we take the per entry maximum to reconstruct the original locality matrix.
2.3.2 AMF loss function
To prevent averaging, rather than extraction of reoccurring subpatterns, we up-weigh overestimation errors, similarly to quantile regression [37], but using a cosine distance instead of mean absolute error. We define the loss function for some locality $i$ as:
@@ 2.3.2 AMF loss function
where $||\,||_{2}$ are $L2$ norms, $\boldsymbol{\hat{C}}_{i}$ is the locality reconstruction (Equation 4), $\odot$ is the Hadamard product, $d_{cos}$ is cosine distance defined as $d_{cos}(\mathbf{x},\mathbf{y})=(1-cos(\mathbf{x},\mathbf{y}))$, MAE is the mean absolute error and $\alpha$ the belonging regularization hyperparameter, and $\mathbf{w}_{i}$ is an overestimation mask, defined as:
@@
where $w_{ic}$ is the weight for locality $i$ and correlation $c$, and $\nu$ is an up-weighing hyperparameter (Appendix A.3). The cosine distance is the main term of the loss function, whose semantics we deem appropriate for module extraction because we are more concerned with patterns of correlations than their absolute values. The scale invariant cosine distance is scaled by a locality’s $L2$ norm to focus on explaining localities showing many high correlations, as opposed to few correlations or mostly weak correlations. The MAE is a regularization term designed to match the scale of the original correlations. This allows for the evaluation of overestimation in the final solutions (Appendix A.3).
2.3.3 Deciding on the number of modules
We decide the number of modules similarly to an approach employed for non-negative matrix factorization (NMF) [38]. For a given number of modules $M$, we run the AMF method $R$ times using different random initializations of matrices $\boldsymbol{M}$ and $\boldsymbol{P}$. To assess the stability of the solution, we find $M$-medoids clusters using all $R\times M$ modules and then calculate the silhouette width based on cosine similarity. A higher silhouette width indicates that similar modules were extracted across the different runs. We repeat this process for various values of $M$, and select a solution with a relatively low number of modules that provides a good trade-off between stability and accurate reconstruction. We argue that lower values of $M$ are desirable, as they will force the optimization to find more global, shared patterns. However, too low a number might force different subpatterns to merge, which should also be reflected in poorer reconstructions. Once a number of modules is chosen, for the final modules, we select the run that has the lowest loss. We comment on limitations in Section 4.

### v2

@@ Abstract
Unsupervised black-box models are drivers of scientific discovery, yet are difficult to interpret, as their output is often a multidimensional embedding rather than a well-defined target. While explainability for supervised learning uncovers how input features contribute to predictions, its unsupervised counterpart should relate input features to the structure of the learned embeddings. However, adaptations of supervised model explainability for unsupervised learning provide either single-sample or dataset-summary explanations, remaining too fine-grained or reductive to be meaningful, and cannot explain embeddings without mapping functions. To bridge this gap, we propose LAVA, a post-hoc model-agnostic method to explain local embedding organization through feature covariation in the original input data. LAVA explanations comprise modules, capturing local subpatterns of input feature correlation that reoccur globally across the embeddings. LAVA delivers stable explanations at a desired level of granularity, revealing domain-relevant patterns such as visual parts of images or disease signals in cellular processes, otherwise missed by existing methods.
Machine Learning, ICML
@@ 1 Introduction
Pragmatically, an explanation can be defined as an answer to a “why” question (Fraassen, 1988). Traditional models, like Newton’s law of gravitation, constitute explanations as they allow reasoning about observable phenomena through their rigorous mathematical description. Similarly, stochastic phenomena that lend themselves well to statistical modeling, such as linear regression, are explainable through probabilistic reasoning. For more complex problems such as facial recognition or protein structure prediction, expert-led modeling is superseded by machine learning (ML), where flexible model architectures and automated end-to-end optimization algorithms take center stage. ML has seen considerable progress and success in leveraging high-dimensional data for scientific discovery across fields like chemistry, mathematics, and biology (Jumper et al., 2021; Davies et al., 2021; Movva et al., 2019; Ching et al., 2018). However, even when ML models make reliable predictions or reflect observed phenomena, reasoning about how they do so is notoriously difficult as most are “black boxes”: the complexity of the data and operations required to produce an output often exceeds what is practically comprehensible (Lipton, 2017; Humphreys, 2009; Beisbart and Räz, 2022). Limited insight into model behavior is a major barrier to adoption of ML for decision making, especially in critical areas such as healthcare (Rudin, 2019; Murdoch et al., 2019; Ching et al., 2018; Wang et al., 2023a). To improve understanding of ML, the community uses either intrinsically interpretable but restrictive models, or explainability methods that allow for post-hoc interpretation of black-box models (Rudin, 2019; Higgins et al., 2018; Schölkopf, 2019; Murdoch et al., 2019; Molnar, 2022).
@@
The popularity and complexity of black-box ML poses challenges to existing explainability methods, creating a need for new solutions. At its core, post-hoc explainability aims to enable reasoning about the output of a model in terms of its treatment of the input data. This allows cross-referencing of model behavior with domain knowledge for verification, assessment of model reliability, and mitigation of unwanted biases. Existing explainability methods cover a wide range of approaches, including feature or sample importance, anchors, and counterfactual explanations, among others (Molnar, 2022). However, most are developed to explain supervised ML models, whose output tends to be a well-defined target label or quantity. This ignores unsupervised black-box ML, designed to reveal latent structure with limited prior knowledge, which can contribute meaningfully to scientific discovery. Indeed, latent embeddings based on autoencoder and transformer architectures have shown effectiveness in various tasks (Brown et al., 2020; Ramesh et al., 2021; Makrodimitris et al., 2024), some serving as a basis of foundation models (Bommasani et al., 2022). Likewise, manifold learning is prominently used in bioinformatics for data exploration, trajectory inference, cell type prediction, spatial domain identification, and multiomics integration (Kobak and Berens, 2019; Brbić et al., 2022; Saelens et al., 2019; Wolf et al., 2019; Haviv et al., 2024).
@@
Crucially, most unsupervised black-box models output multi-dimensional latent embeddings. Relating latent dimensions to the input data is less straightforward, as the axes of the learned latent space do not correspond to known quantities and are therefore difficult to interpret (Crabbé and Schaar, 2022). Disentangled representations have shown mixed results, suggesting that latent dimensions cannot generally be explained independently from one another (Crabbé and Schaar, 2022; Locatello et al., 2020). As such, a promising direction for explainability of unsupervised black-box models is to focus on the relation between the relative spatial organization of the embeddings and the underlying data. Moreover, many manifold learning methods only produce latent representations of the samples, without a parametrized mapping function to embed new data (Murphy, 2022). This means that all information is contained within the embeddings, making the link between latent organization and underlying data crucial for reasoning about the model.
@@
Literature on explainability of unsupervised black-box latent embeddings is limited. Most notably, Crabbé and Schaar (2022) propose a framework adapting feature and sample importance explainability for supervised models to the unsupervised setting. This contribution importantly enables explanations for unsupervised ML using a wide range of methods. However, these yield either single-sample or dataset-summary explanations, which ignore the organization of latent embeddings and are typically too fine-grained or too global to meaningfully understand the underlying model. This is especially relevant as many black-box models are non-linear and seek to capture local structure in the original high-dimensional data. In fact, the popular t-SNE and UMAP manifold learning algorithms  (Meilă and Zhang, 2024; Maaten and Hinton, 2008; McInnes et al., 2020) explicitly aim to preserve local sample neighborhoods of the input in a lower-dimensional latent space. We therefore argue that effective explainability of latent embeddings requires some strategy of relating samples to one another based on their organization in the latent space. While manifold learning methods have been extensively studied in terms of neighborhood retention (Han et al., 2022) and hyperparameter effects (Becht et al., 2019; Kobak and Berens, 2019; Wang et al., 2021; Xia et al., 2024), no research has focused on explaining the organization of the embeddings as a function of the input data. More manual approaches relate spatial organization to original features (Liu et al., 2019), but relying purely on prior knowledge, metadata, or target features of interest, and are infeasible for high-dimensional data.
Here, we propose Locality-Aware Variable Associations (LAVA), a post-hoc model-agnostic method to explain learned latent embeddings by relating input features to latent spatial organization. LAVA generates explanations as modules capturing local subpatterns of input feature correlation reoccurring globally in representative neighborhoods throughout the latent space. LAVA modules therefore describe latent structure at a desired level of granularity, in terms of input feature covariation in the underlying data. We use LAVA to explain embeddings of the MNIST and single-cell kidney datasets, showcasing its ability to identify visually and biologically relevant modules of pixel and gene expression correlation, respectively. We further characterize LAVA explanations and assess their stability. Finally, we also compare LAVA to alternative explainability and module extraction methods.
Figure 1: LAVA method overview. a, LAVA explains embeddings of a dataset (example UMAP embeddings of MNIST). b, Based on neighborhood size $n$, LAVA represents embeddings as a set of localities: overlapping neighborhoods centered around probes. c, For each locality, LAVA calculates feature-to-feature correlations using the original samples of that locality (top right, locality correlations; bottom right, heatmap showing the sum of correlations per pixel, with added lines connecting correlated pixels). d, LAVA extracts modules, shared subpatterns of correlation extracted from localities that serve as explanations, describing feature covariation across the embeddings.
@@ 2 Methodology
LAVA creates explanations relating local organization of latent embeddings to feature covariation of the underlying data in three steps: (1) locality definition, (2) locality representation, and (3) module extraction (Figure 1). Step 1 defines the scale and reference points (probes) of the neighborhoods, or localities, to use for capturing local organization. Step 2 creates a representation of each locality, linked to the original features via their correlation. Step 3 produces the actual explanations, consisting of modules of common correlation subpatterns across the localities.
@@ 2.1 Locality Definition
To reflect the local organization of latent embeddings, LAVA uses a smaller number of localities, centered around probe points (not samples) spread across the latent space (Figure 1b). A locality centered at probe $p$ is an $n$-sized Euclidean neighborhood containing the $n$ samples nearest to $p$ in the latent space. The number of localities $\ell\leq E$ is given by $\ell=E\times o/n$, based on the number of embeddings $E$, neighborhood size $n$, and factor $o$ controlling the extent of overlap between localities (Appendix A.1).
Locality placement. The goal of locality placement is to determine the locations of the $\ell$ probes around which the $\ell$ localities will be formed. To ensure that the localities are representative of the latent embedding structure, LAVA optimizes locality in-degree centrality towards neighborhood in-degree centrality (Figures S12, S18, and S28). In other words, the proportion of probe-centered localities including a given sample is made to approximate the proportion of all possible sample-centered $n$-neighborhoods including that same sample. Practically, LAVA finds the locations of the $\ell$ probes as cluster centers determined using a custom weighted $k$-means formulation, where $k=\ell$. The clustering uses the following sample weighting scheme:
$\centering w_{i}=\left(\frac{in\_neighbor_{i}}{E}\right)^{\alpha}\times\left(\frac{1}{\bar{d}_{i}^{n}}\right)^{\beta},\@add@centering$ (1)
where $w_{i}$ is the weight of sample $i$, $in\_neighbor_{i}$ is the in-degree centrality of sample $i$ using sample-based $n$-neighborhoods, $\bar{d}_{i}^{n}$ is the average distance from sample $i$ to its $n$ nearest neighbors, and $\alpha$ and $\beta$ are the parameters optimized using the DIRECT algorithm (MacQueen, 1967; Jones et al., 1993). We define the cost function for DIRECT as the Manhattan distance between the neighborhood and locality in-degree centralities of each sample, as follows:
$\centering\mathcal{L}(\alpha,\beta)=\sum_{i=1}^{E}\left|\frac{in\_neighbor_{i}}{E}-\frac{in\_locality_{i,\alpha,\beta}}{\ell}\right|,\@add@centering$ (2)
where $in\_locality_{i,\alpha,\beta}$ is the number of localities around the resulting $\ell$ cluster centers that include $i$ as a neighbor. Details and pseudocode in Appendix A.1.
@@ 2.2 Locality Representation
After defining localities, LAVA links these local organizations of embeddings to the original data. This is achieved by representing each locality through pairwise associations of the original features, calculated using only the samples within that locality. In this version of LAVA, association is defined as the absolute Spearman correlation, capturing the strength of monotone association between features as values in $[0,1]$ (Figure 1c). Therefore, each locality is represented by a vector of $D\choose 2$ pairwise correlations, where $D$ is the number of features. Unlike Pearson correlations, Spearman correlations are not scale dependent and thus more flexible in the type of association they capture, while remaining computationally inexpensive and interpretable downstream. Individual correlations are filtered out (set to 0) if one of the original features has the same value in $75\%$ of the neighborhood (Appendix A.2).
Figure 2: LAVA association matrix factorization (AMF). a, Locality $i$ is reconstructed by multiplying each module by its presence for $i$, then taking the maximum per entry across the $M$ extracted modules. b, Localities are represented as matrix $\boldsymbol{C}$: each row is a locality, each column a pair of features, and each entry the pairwise feature correlation in that locality. Module extraction approximates $\boldsymbol{C}$ with a maximum per entry based on modules $\boldsymbol{M}$ and their presences $\boldsymbol{P}$: it calculates an outer product between each column of $\boldsymbol{P}$ (superscript) and each row of $\boldsymbol{M}$, yielding a tensor of $M$ matrix slices of the same shape as $\boldsymbol{C}$ (Equation 5). Each slice represents module-contributed correlations across the localities. The $M$ dimension reduced by taking the per element maximum yields the reconstructed locality dataset.
@@ 2.3 Module Extraction
Finally, LAVA extracts modules of reoccurring feature correlation subpatterns across localities, as explanations of the latent embeddings (Figure 2). Modules adhere to two properties: (1) modules capture reoccurring rather than average patterns across localities, as modules based on averages of localities would not be accurate representations of correlations observed at each locality individually; (2) modules can be co-present in the same locality, allowing decomposition of the correlation patterns.
Modules are identified based on the matrix of locality representations $\boldsymbol{C}$, with rows denoting the $\ell$ localities and columns representing the $D\choose 2$} pairwise correlations. To identify modules, we propose the Association Matrix Factorization (AMF) method, which factorizes a matrix like $\boldsymbol{C}$ into two module matrix $\boldsymbol{M}$ and presence matrix $\boldsymbol{P}$. Module matrix $\boldsymbol{M}$ is of size $M\times{D\choose 2}$, where each row represents one of $M$ modules, columns denote correlations, and each row vector expresses a subpattern of correlations of a module. Presence matrix $\boldsymbol{P}$ is of size $\ell\times M$, with each entry quantifying the presence of one of the $M$ modules in one of the $\ell$ localities. Presence and module matrix entries are restricted to real values in $[0,1]$, and reconstruct locality correlations using the formulation below.
AMF formulation. The formulation of AMF enables modules to overlap across localities. Crucially, as correlations or measures of association more broadly cannot be considered additive, we avoid summing correlations and instead use a per element maximum function. This allows modules to contribute overlapping subpatterns of correlations without direct interaction in the reconstruction. To that effect, we reconstruct each correlation $j$ of each locality $i$ by taking the maximum of the modules’ corresponding correlation values multiplied by the presences of those modules at locality $i$:
@@
where $\odot$ is the Hadamard product and single letter subscripts indicate matrix rows. The reconstruction of a locality $i$ can then be written as a per element maximum across products of the modules and their presences in locality $i$ (Figure 2a):
$\centering\boldsymbol{\hat{C}}_{i}=\max\limits_{1\leq j\leq{D\choose 2}}\left(\boldsymbol{P}_{i1}\times\boldsymbol{M}_{1}^{T},...,\boldsymbol{P}_{iM}\times\boldsymbol{M}_{M}^{T}\right),\@add@centering$ (4)
@@
$\centering\boldsymbol{P}^{m}\otimes\boldsymbol{M}_{m}^{T}=\begin{bmatrix}\boldsymbol{P}_{1m}\boldsymbol{M}_{m1}&\ldots&\boldsymbol{P}_{1m}\boldsymbol{M}_{m{D\choose 2}}\\ \vdots&\ddots&\vdots\\ \boldsymbol{P}_{\ell m}\boldsymbol{M}_{m1}&\ldots&\boldsymbol{P}_{\ell m}\boldsymbol{M}_{m{D\choose 2}}\end{bmatrix}$,\@add@centering (5)
where the superscript $m$ denotes the column of the matrix $P$ (the presence vector of some module $m$) and $\otimes$ is the outer product. Intuitively, the reconstruction of all localities is related to matrix multiplication of $P$ and $M$, but instead of summing up entries across $M$, we take the maximum. In tensor terminology, outer products from Equation 5 for each of the modules produce $M$ slices of size $\ell\times{D\choose 2}$, over which we take the per entry maximum to reconstruct $C$.
AMF loss function. We define the loss for locality $i$ as:
@@
where $||\,||_{2}$ are $L2$ norms, $\boldsymbol{\hat{C}}_{i}$ is the locality reconstruction (Equation 4), $\odot$ is the Hadamard product, $d_{\cos}$ is cosine distance defined as $d_{\cos}(\mathbf{x},\mathbf{y})=(1-\cos(\mathbf{x},\mathbf{y}))$, MAE is the mean absolute error with respective regularization hyperparameter $\gamma$. Overestimation mask $\mathbf{w}_{i}$ is defined as:
@@
where $w_{ic}$ is the weight for locality $i$ and correlation $c$, and $\nu$ is an up-weighing hyperparameter (Appendix A.3).
Figure 3: LAVA stability experiments. Top results for MNIST, bottom for KPMP UMAP embeddings. a, Cosine similarity of average correlations of locality placements when differing the overlap hyperparameter. b, Statistics of module extraction runs. c-d, Cosine similarity between presence-weighted averaged modules c, across different locality placements and d, within one locality placement.
The cosine distance is the main term of the loss function, whose semantics we deem appropriate for module extraction because we are more concerned with patterns of correlations than their absolute values. However, to prevent extraction of average, rather than reoccurring patterns, we up-weigh overestimation errors, similarly to quantile regression (Koenker and Hallock, 2001) (Section 3.2). The MAE ensures modules are on the same scale as the locality correlations, despite the scale invariant cosine distance. The loss is optimized with ADAM (Kingma and Ba, 2017) (Appendix A.3).
Number of modules. To decide the number of modules, we follow an approach used for non-negative matrix factorization (NMF) (Alexandrov et al., 2013), based on reconstruction and stability. For a range of values of $M$, we run AMF $R$ times using different random initializations of matrices $\boldsymbol{M}$ and $\boldsymbol{P}$. To assess stability, we find $M$-medoids clusters using all $R\times M$ modules and calculate silhouette width (Rousseeuw, 1987) based on cosine similarity. Higher stability indicates similar modules extracted across runs. Lower values of $M$ are desirable to guide the optimization towards more global shared patterns. However, too low a number might force subpatterns to merge, which should be reflected in poorer reconstruction. We thus look for an elbow of the loss across $M$ values, resolving ambiguity by favoring higher stability (Section 3.1).